a 90 inch pipe is cut into two pieces. one piece is 20 in more than the other. find the lenth of each piece.
Answer by orca(409)

Let x be the shorter piece, the the longer one is x + 20.
As the total length is 90, we have
x + x+20 = 90
Solving for x, we have
2x + 20 = 90
2x = 70
x = 35
So the shorter one is 35 inches, and the longer one is x + 20 = 35 + 20 = 55 inches.